Can you please help me, I have posted two of four questions below. I was hoping you could help me understand the first two so that I could do the second two without help. I am very confused.
For questions 16-20, classify the given numbers as real and rational, real and irrational, or complex.
16. (2 points) (4)^1/2 + 2
17. (2 points) 6+0i

Answer by richwmiller(17219)

16. 4^(1/2)+2 is sqrt(4)+2 =-2+2 =0 and 2+2=4
0 and 4
Both are real
17. 6 +0i
The form is complex but since the coefficient is zero, it is real.